A company has 7 warehouses, only 4 of which have a particular product in stock. A salesperson calls the warehouses in a random order until a warehouse with the product is reached. Let the random variable X be the number of calls made by the salesperson. What is the P(X=2)?

Round your answer to the nearest hundredth.
Answer #1

Ans:

Use Geometric distribution:

Probability of success,p=4/7

P(X=k)=(1-p)k-1*p

P(X=2)=(1-(4/7))2-1*(4/7)

=(3/7)*(4/7)=12/49=0.24